Hi okay, I dont understand the question "if two angles are complementary and the larger angle is 5 less than 3 times the smaller, then find these two angle measurements. Draw and Label a sketch to help you" I drew a sketch and labeled it but it did not help me, I still dont get it! Please Help Me! Thanks very much!
Found 2 solutions by rfer, mananth:
Answer by rfer(16322)

x+3x-5=90
4x=95
x=23.75
3x-5=66.25

Answer by mananth(16949)

let the angle be x
the complementary will be 90 -x
..
x= 3(90-x)-5
x= 270-3x-5
x=265-3x
add 3x
x+3x=265-3x+3x
4x=265
x=66.5 deg.one angle
..
90-66.5 = 23.5 deg is the other angle
..
m.ananth@hotmail.ca